# Tunneling in multi-site mesoscopic quantum Hall circuits

### v1

@@ Abstract
Transport properties of the single- and two-site mesoscoipc quantum Hall (QH) circuits at high transparencies can be described in terms of the lowest-order backscattering perturbations, and mapping to the boundary sine-Gordon model can be exploited in full generality. While the higher-order backscattering processes are exactly marginal in the case of corresponding three-site circuits, they become crucial in a device with four or more sites. Here, we explore the transport properties of a multi-site QH circuit with special focus on that with four sites, and report their unique quantum critical behaviors that can be accessed via transport measurements. Tunneling phenomena in multichannel QH circuits based on multi-site geometry are also investigated, and a promising route to realizing different aspects of quantum critical phenomena is offered.
@@ I Introduction
In a typical low-dimensional quantum system, the interplay among strong electron interactions, quantum interferences, and different degeneracies generally gives rise to a number strongly correlated phenomena [1]. Manifestations of these effects can also be studied using a well-controllable mesoscopic device, called the single-electron transistor [2]. It can be realized as a mesoscopic quantum Hall (QH) circuit consisting of a floating metal grain coupled to one-dimensional QH edge channels [3, 4, 5]. This metal-semiconductor hybrid structure can exhibit various quantum effects such as Coulomb blockade, Kondo effect, charge and spin fractionalization, Luttinger liquid behavior and quantum phase transition [3, 6, 7, 8, 9, 10, 11, 12, 13]. These effects constitute a paradigm in the field of strongly correlated electron physics.
@@
A number of remarkable phenomena have been predicted to emerge in a two-site QH circuit comprising coupled hybrid metal-semiconductor islands due to the competition between the screening of each individual grain charge and the mediated charge coupling between the two grains [14, 15], i.e., interplay of local and nonlocal correlation effects (see also Refs. [16, 17, 18, 19]). Such competition also results in local parafermionic excitation [20]. We note that, unlike the Majoranas (that could be realized in a single-site QH circuit [21, 8]), parafermions are much more complex, since they cannot arise in non-interacting models, being the result of strong correlation effects [22]. Therefore, the recent experiment [14] and theoretical predictions [15, 20] open coveted pathways to further understand and manipulate the exotic parafermionic state of matter using a highly-tunable nanoelectronic device without topological orders, and also to realizing different aspects of zero-temperature quantum critical phenomena [23].
@@
One of the promising research directions is then to scale up such hybrid structures to form a lattice that could provide different routes to constructing more complex Fibonacci anyons by leveraging the exotic features of generalized coupled-impurities models. These multi-site QH circuits can also provide a useful way of understanding the effects of disorder on QH edges due to QH puddle formation [24] (see also Ref. [25]). In addition, they show some novel aspects of heat transport properties [26, 27, 28, 29, 30, 31]. It is important to note that owing to having infinitely large numbers of electrons in the metal islands, each grain in the chain, essentially behaves identical to each other. The well-defined QH edge states implemented via quantum point contacts in the metal-semiconductor hybrid structures also provide unprecedented tuning of the device. Therefore, multi-site QH circuits could be used to simulate different model quantum systems of strongly correlated physics, i.e. they could capture the essential features of a real material.
@@
It is well-understood that the combined effects of a dynamical Coulomb blockade and integer QH effect generate a form of a Luttinger liquid (LL) [32]. A single channel QH circuit consisting of $N$ grains can be described by an equivalent model of an impurity in LL with the Luttinger parameter $K=1/(N+1)$, or a fractional QH effect with filling fraction $\nu=K$ [15]. As long as there are no tunneling/backscattering events in the system, all the properties of multi-site QH circuits can be achieved just by the usual circuit analysis [33].
@@
Figure 1: Schematic of the four-site quantum Hall circuit consisting of four essentially identical metallic islands connected by five external quantum point contacts (QPCs). Each island is connected to separate external gates (green metallic plates). The red dots represent the backscattering centers of fully-tunable QPCs. The right-moving edge state on the left is voltage $V$ biased, and charge current is measured at the QPC5. The symbol $\Phi_{\rm jR/L}$ stands for the bosonic field describing the right/left moving edge state.
@@
In the case of a finite backscattering between counter-propagating edge states forming the circuit, the nature of relevant perturbations depends on the typical value of $\nu$. In the standard problem of tunneling through a weak barrier represented by a potential scatterer $\mathcal{V}(x)$ localized in the vicinity of $x=0$, the relevant boundary perturbations correspond to the process of backscattering of $n$ electrons from one Fermi point to the other by momenta $2k_{\rm F}$ [34, 35]. These perturbations are also called $\mathcal{V}(2nk_{\rm F}),\;n=1,2,\cdots$, scattering processes. For all values of $\nu$, the $\mathcal{V}(2k_{\rm F})$ scattering is always relevant in the renormalization group (RG) sense, and all higher order scatterings that satisfy $n^{2}\nu<1$ are also relevant perturbation [34]. This shows that only the single-site $N=1$, and two-site $N=2$ QH circuits can fully be described by the lowest order perturbation $\mathcal{V}(2k_{\rm F})$ where mapping to the boundary sine-Gordon model can be exploited in full generality [36, 37]. For $N=3$, the $\mathcal{V}(4k_{\rm F})$ process are exactly marginal, and just introduce trivial renormalization effects. Therefore, to study the transport properties of multi-site QH circuits at large transparencies, it is crucial to also consider the higher order perturbation for $N\geq 4$. The same thing is also true even for $N=1$ and $N=2$ but with the fractional QH edges described by Laughlin state with filling factor $\nu=1/m$ for $m$ being an odd integer [38].
This work explores the transport properties of a four-site QH circuit 1 as a prototypical setup where higher order backscattering process are the relevant perturbations (increasing more sites does not introduce fundamentally new effects). We identify the associated zero temperature quantum critical point by uncovering the parameter space in gate voltages and contact resistances at which all the relevant backscattering processes interfere destructively, giving maximal conductance irrespective of temperature and/or voltage. We also evaluate the differential conductance of the device at two different limits, and discuss their scaling behaviors in the vicinity of quantum critical point. We introduce tunneling phenomena in the multichannel, multi-site mesoscopic QH circuit, and analyze their quantum critical behaviors. An experimentally accessible route to realize numerous singular phenomena in strongly correlated system is proposed. Additionally, the effects of Joules heating on the explored devices are uncovered.
@@
This paper is organized as follows. In Sec. II we present our bosonization-based model to describe the multi-site QH circuits. An exact solution of the model in the absence of backscattering events is given in Sec. III. Section IV is devoted to the construction of an effective low-energy Hamiltonian that takes into account the different tunneling process in the considered system. Quantum critical properties of a multi-site QH circuit are discussed in Sec. V, and their transport properties in different geometries are presented in section VI. In section VII, we introduce the multichannel, multi-site mesoscopic QH circuits and analyze their quantum critical behaviors. A brief discussion about the Joules heating effects in the explored devices is given in Sec. VIII. Finally, we conclude in Sec. IX. Minor mathematical details of our calculations are given in the Appendices.
@@ II Model
The schematic of the experimental setup, which consists of four hybrid metal-semiconductor islands, is shown in Fig. 1. Each island hosts a macroscopically-large number of charge states that can be tuned by applying voltage on the corresponding nearby gates (green metallic plates in the figure). The islands are connected to each other and also to the metallic leads by fully tunable quantum point contacts (QPCs) represented by the red dot signs in the figure. The free electrons in the QPCs can be modeled as pairs of counter-propagating QH edges which are partially covered by the metallic islands. Electron propagation through these chiral edge states can conveniently be studied by bosonizing fermion operators [39]. In this representation, the spinless electrons in five QPCs can be described by the quadratic Hamiltonian of the form [40]
@@
where $\alpha$ is the QPCs number index, $v_{\rm F}$ stands for the Fermi velocity. $\Phi_{\alpha,{\rm L/R}}(x,t)$ represents the bosonic field corresponding to the left/right moving chiral fermions, and satisfies the commutation relation $\left[\Phi_{\rm\alpha,R/L}(t),\Phi_{\rm\alpha,R/L}(t_{1})\right]=\pm i\pi{\rm sign}(t-t_{1})$. Notice that throughout the paper we set $e=\hbar=k_{\rm B}=1$.
The Coulomb interactions in the system of four islands having the same charging energy $E_{C}$ can be accounted for by the constant interaction model [8]
@@
where $j$ is the islands number index, and the symbol $N_{j,\rm g}$ stands for the gate voltage applied to the $j$th grain. The electron number operator of the $j$th island is given, in terms of the bosonic fields defined at $x=0$, by
$\hat{N}_{j}(t)=\frac{1}{2\pi}\left[\Phi_{j,\rm L}-\Phi_{j,\rm R}+\Phi_{j+1,\rm R}-\Phi_{\rm j+1,\rm L}\right].$ (3)
We note that in the setup 1, there is indeed an appreciable amount of inter-island Coulomb interactions. The cross capacitance facilitating the charge-charge coupling on the nearest islands, however, does not allow the DC current to follow. Therefore, as long as we are interested in low-energy charge transport, such charge-charge couplings can be safely neglected as far as we are far from the regime of possible phase transition [41].
The interaction Hamiltonian (2) is quadratic in bosonic fields, and thus can be solved exactly. In addition, if the system can fully be described by equations (1) and (2), the gate voltages can be gouged out as the system looses charge granularity. Small but finite backscattering events mediated by QPCs principally change this picture. Such events result in charge granularity and introduce strongly non-linear boundary interactions of the form $H_{\rm b}=\sum_{\alpha}\sum_{n}\mathcal{V}(2nk_{F})\cos\left[n\left(\Phi_{\alpha,\rm R}-\Phi_{\alpha,\rm L}\right)\right]$. For the low-energy description of the setup comprising of $N$ islands, the QPC index $\alpha$ takes the values $\alpha=1,2,\cdots,N+1$, and the integer $n=1,2,\cdots\lfloor\sqrt{N+1}\rfloor$. Therefore, the backscattering in the QPCs of the device (1), i.e., $N=4$, is accounted for by
$H_{\rm b}$ $=\sum_{\alpha=1}^{5}\Big[U_{\alpha}\cos\left(\Phi_{\alpha,\rm R}-\Phi_{\alpha,\rm L}\right)$
$\qquad\qquad+V_{\alpha}\cos 2\left(\Phi_{\alpha,\rm R}-\Phi_{\alpha,\rm L}\right)\Big],$ (4)
where $U_{\alpha}\propto\mathcal{V}(2k_{F})$, $V_{\alpha}\propto\mathcal{V}(4k_{F})$ and they depend on the band cutoff $D$ in a linear and a quartic manner respectively. In the following, we solve the interaction Hamiltonian (2) exactly, and take into account the non-linearities (4) perturbatively in small backscattering coefficients [8].
@@ III Exact account of the interactions
In the absence of backscattering events, the continuous flow of chiral fermions along the QH edges described by the Hamiltonian $H_{\rm p}=H_{0}+H_{C}$ represents the plasmonic excitations traveling along the chiral edges. The latter can be accounted for by the standard equation of motion approach [33, 42, 43]. This allows one to express all the bosonic fields $\Phi_{\alpha,\rm R/L}(x,\omega)$ in terms of corresponding free fields $\Phi^{0}_{\alpha,\rm R/L}(x,\omega)$ satisfying the well-known commutation relations and have standard correlators
@@
To study the transport properties, it is more convenient to introduce the compact bosonic fields $\Phi_{\alpha}\equiv\Phi_{\alpha,\rm R}-\Phi_{\alpha,\rm L}$, and corresponding free fields $\Phi^{0}_{\alpha}$. It is straightforward to show, by solving the equation of motion for the Hamiltonian $H^{\prime}\equiv H_{0}+H_{C}$, that the matrix $\Phi=\left(\Phi_{1},\Phi_{2},\Phi_{3},\Phi_{4},\Phi_{5}\right)^{\top}$ satisfies
@@
for some matrix $\mathbb{A}$ depending on the interaction strength and number of edge channels. The matrix $\mathbb{N}$ accounting for the gate voltages takes the form
@@
From the equations (25) and (26), it is seen that out of five chiral bosonic modes, only one mode, $\Phi_{a}$ remains gapless, and the other four are gapped by large charging energy. For the low-energy description of the problem, the gapped modes can be integrated out, and thus the dynamics are governed solely by the free mode $\Phi_{a}$. The equation (26) also implies that the scaling dimensions of the boundary perturbations expressed by the first and second terms of the equation (4) are $1/5$ and $4/5$ respectively, and thus both are relevant perturbations in RG sense.

### v2

@@ Abstract
Transport properties of single- and two-site mesoscopic quantum Hall (QH) circuits at high transparencies can be described in terms of the lowest-order backscattering processes, enabling a mapping to the boundary sine-Gordon model. We show that this description breaks down in circuits with four or more sites, where higher-order backscattering processes become relevant and qualitatively modify the low-energy physics, while remaining exactly marginal in three-site geometries. Focusing on the four-site circuit, we derive an effective low-energy theory that captures the resulting interaction-driven physics and reveal the emergence of unique quantum-critical points. In the vicinity of these critical points, we obtain universal conductance and scaling behavior and establish the robustness of the associated non-Fermi liquid physics. We further introduce tunneling in multichannel multi-site QH circuits and propose a promising route for realizing diverse quantum-critical phenomena. We show that a boundary sine-Gordon description can be restored in multichannel multi-site QH circuits by appropriately looping selected edge channels, a procedure that is experimentally feasible. Finally, we analyze the non-equilibrium heating effects relevant to transport measurements in QH circuits. Altogether, our results establish multi-site QH circuits as a versatile and highly controllable platform for simulating interaction-driven quantum critical phenomena.
@@ I Introduction
In a typical low-dimensional quantum system, the interplay among strong electron interactions, quantum interference, and different degeneracies generally gives rise to a number of strongly correlated phenomena [1]. Manifestations of these effects can also be studied using a well-controllable mesoscopic device, called the single-electron transistor [2]. It can be realized as a mesoscopic quantum Hall (QH) circuit consisting of a floating metal grain coupled to one-dimensional QH edge channels [3, 4, 5]. This metal-semiconductor hybrid structure can exhibit various quantum effects such as Coulomb blockade, the Kondo effect, charge and spin fractionalization, Luttinger liquid behavior, and quantum phase transitions [3, 6, 7, 8, 9, 10, 11, 12, 13]. These effects constitute a paradigm in the field of strongly correlated electron physics.
@@
A number of remarkable phenomena have been predicted to emerge in a two-site QH circuit comprising coupled hybrid metal-semiconductor islands due to the competition between the screening of each individual grain charge and the mediated charge coupling between the two grains [14, 15], i.e., the interplay of local and nonlocal correlation effects (see also Refs. [16, 17, 18, 19]). Such competition also results in local parafermionic excitations [20]. We note that, unlike the Majorana modes (which could be realized in a single-site QH circuit [21, 8]), parafermions are much more complex, since they cannot arise in noninteracting models, being the result of strong correlation effects [22]. Therefore, the recent experiment [14] and theoretical predictions [15, 20] open up coveted pathways to further understand and manipulate the exotic parafermionic state of matter using a highly tunable nanoelectronic device without topological order. In addition, by adjusting the number of channels connecting the two sites, exotic zero-temperature quantum critical phenomena with diverse critical exponents $\eta\leq 1/2$ can be realized [23].
@@
One of the promising research directions is to scale up such hybrid structures to form a lattice that could provide different routes to constructing more complex Fibonacci anyons by leveraging the exotic features of generalized coupled-impurity models. These multi-site QH circuits can also provide a useful way of understanding the effects of disorder on QH edges due to QH puddle formation [24] (see also Refs. [25, 26]). In addition, they show novel aspects of heat transport [27, 28, 29, 30, 31, 32, 33]. It is important to note that owing to the presence of an infinitely large number of electrons in the metal islands, each grain in the chain, essentially behaves identically to the others. The well-defined QH edge states implemented via quantum point contacts in the metal-semiconductor hybrid structures also provide unprecedented tuning of the device. Therefore, multi-site QH circuits could be used to simulate different model quantum systems in strongly correlated physics, i.e., they could capture the essential features of a real material [34].
@@
It is well understood that the combined effects of a dynamical Coulomb blockade and the integer QH effect generate a form of Luttinger liquid (LL) [35]. A single channel QH circuit consisting of $N$ grains can be described by an equivalent model of an impurity in an LL with the Luttinger parameter $K=1/(N+1)$, or of a fractional QH effect with filling fraction $\nu=K$ [15]. As long as there are no tunneling/backscattering events in the system, all the properties of multi-site QH circuits can be achieved simply through the usual circuit analysis [36].
@@
Figure 1: Schematic of the four-site quantum Hall circuit consisting of four essentially identical metallic islands connected by five external quantum point contacts (QPCs). Each island is connected to a separate external gate (green metallic plates). The red dots represent the backscattering centers of fully tunable QPCs. The right-moving edge state on the left is biased by a voltage $V$, and the charge current is measured at QPC5. The symbol $\Phi_{\rm jR/L}$ stands for the bosonic field describing the right-/left-moving edge state.
@@
In the case of a finite backscattering between counter-propagating edge states forming the circuit, the nature of the relevant perturbations depends on the typical value of $\nu$. In the standard problem of tunneling through a weak barrier represented by a potential scatterer $\mathcal{V}(x)$ localized in the vicinity of $x=0$, the relevant boundary perturbations correspond to the process of backscattering of $n$ electrons from one Fermi point to the other by momenta $2k_{\rm F}$ [37, 38]. These perturbations are represented by the Fourier components of the backscattering potential, i.e., $\mathcal{V}(2nk_{\rm F}),\;n=1,2,\cdots$, and are commonly called as $n$th-order backscattering processes. For all values of $\nu$, the $\mathcal{V}(2k_{\rm F})$ scattering is always relevant in the renormalization group (RG) sense, and all higher order scatterings that satisfy $n^{2}\nu<1$ are also relevant perturbation [37]. This shows that only the single-site $N=1$, and two-site $N=2$ QH circuits can fully be described by the lowest-order backscattering processes $\mathcal{V}(2k_{\rm F})$ where mapping to the boundary sine-Gordon model can be exploited in full generality [39, 40]. For $N=3$, the $\mathcal{V}(4k_{\rm F})$ processes is exactly marginal, and introduce only trivial renormalization effects. Therefore, to study the transport properties of multi-site ($N\geq 4$) QH circuits at large transparencies, it is crucial to also consider the higher-order backscattering processes. The same thing is also true even for $N=1$ and $N=2$ but with the fractional QH edges described by the Laughlin state [41] with filling factor $\nu=1/m$, where $m$ is an odd integer.
In the present work, we broadly focus on two directions. First, we explore the transport properties of a four-site QH circuit, shown in Fig. 1, which serves as a prototypical setup where higher-order backscattering processes constitute the relevant perturbations; increasing the number of sites does not introduce fundamentally new effects. Second, we introduce and analyze tunneling phenomena in multichannel realizations of multi-site mesoscopic QH circuits. Together, we establish multi-site QH circuits as a versatile and experimentally accessible platform for realizing and probing quantum criticality, interference-driven transport, and non-Fermi-liquid behavior in strongly correlated mesoscopic systems.
@@
This paper is organized as follows. In Sec. II we present our bosonization-based model to describe the multi-site QH circuits. An exact solution of the model in the absence of backscattering events is given in Sec. III. Section IV is devoted to the construction of an effective low-energy Hamiltonian that takes into account the different tunneling processes in the considered system. Quantum critical properties of a multi-site QH circuit are discussed in Sec. V, and their transport properties in different geometries are presented in sections. VI and  VII. In sections VIII and  IX, we introduce the multichannel, multi-site mesoscopic QH circuits and analyze their quantum critical behaviors. Discussion of the Joule heating effects in the explored devices is given in Sec. X. Finally, we conclude in Sec. XI. Minor mathematical details of our calculations are given in the Appendices.
@@ II Model
The schematic of the experimental setup, consisting of four hybrid metal-semiconductor islands, is shown in Fig. 1. Each island hosts a macroscopically large number of charge states, which can be tuned by applying voltages to the corresponding nearby gates (green metallic plates in the figure). The islands are connected to each other, as well as to the metallic leads by fully tunable quantum point contacts (QPCs) represented by the red dots in the figure. The free electrons in the QPCs can be modeled as pairs of counter-propagating QH edge states that are partially covered by the metallic islands. Electron propagation through these chiral edge states can conveniently be studied using bosonization of fermion operators [42]. In this representation, the spinless electrons in the five QPCs are described by the quadratic Hamiltonian [43]
@@
where $\alpha$ labels the QPCs and $v_{\rm F}$ denotes the Fermi velocity. $\Phi_{\alpha,{\rm L/R}}(x,t)$ represents the bosonic field corresponding to the left/right moving chiral fermions, and satisfies the commutation relation $\left[\Phi_{\rm\alpha,R/L}(t),\Phi_{\rm\alpha,R/L}(t_{1})\right]=\pm i\pi{\rm sign}(t-t_{1})$. Throughout this paper we set $e=\hbar=k_{\rm B}=1$.
The Coulomb interactions in the system of four islands, each with same charging energy $E_{C}$, are described by the constant-interaction model [8]
@@
where $j$ labels the islands, and $N_{j,\rm g}$ denotes a dimensionless parameter proportional to the gate voltage $V_{j,{\rm g}}$, i.e., $eN_{j,{\rm g}}=C_{j,{\rm g}}V_{j,{\rm g}}$, where $C_{j,{\rm g}}$ is the capacitance between the $j$th island and the corresponding gate. In the following, we will use the term “dimensionless gate voltages” or simply “gate voltages” to refer to $N_{j,{\rm g}}$. The electron number operator of the $j$th island is given, in terms of the bosonic fields evaluated at $x=0$, by
$\hat{N}_{j}(t)=\frac{1}{2\pi}\left[\Phi_{j,\rm L}-\Phi_{j,\rm R}+\Phi_{j+1,\rm R}-\Phi_{\rm j+1,\rm L}\right],$ (3)
where we used the abbreviation $\Phi_{j,\rm L/R}(x=0,t)\equiv\Phi_{j,\rm L/R}$. We note that in setup 1, there is indeed an appreciable amount of inter-island Coulomb interaction. The cross capacitance that facilitates the charge-charge coupling between neighboring islands, however, does not allow the dc current to flow. Therefore, as long as we are interested in low-energy charge transport and remain far from a possible phase-transition, such charge-charge couplings can be safely neglected [44].
The interaction Hamiltonian (2) is quadratic in the bosonic fields and can therefore be solved exactly. Moreover, if the system is fully described by equations (1) and (2), the gate voltages can be gauged out, resulting in a system without charge granularity. Small but finite backscattering events mediated by the QPCs fundamentally alter this picture. Such events restore charge granularity and introduce strongly nonlinear boundary interactions of the form $H_{\rm b}=\sum_{\alpha}\sum_{n}\mathcal{V}(2nk_{F})\cos\left[n\left(\Phi_{\alpha,\rm R}-\Phi_{\alpha,\rm L}\right)\right]$. For the low-energy description of a setup comprising $N$ islands, the QPC index $\alpha$ runs from 1 to $N+1$, and the integer $n$ takes the values $n=1,2,\cdots\lfloor\sqrt{N+1}\rfloor$. Therefore, the backscattering in the QPCs of device (1), i.e., $N=4$, is described by
$H_{\rm b}(t)$ $=\sum_{\alpha=1}^{5}\Big[U_{\alpha}\cos\left[\Phi_{\alpha,\rm R}(0,t)-\Phi_{\alpha,\rm L}(0,t)\right]$
$\qquad+V_{\alpha}\cos 2\left[\Phi_{\alpha,\rm R}(0,t)-\Phi_{\alpha,\rm L}(0,t)\right]\Big],$ (4)
where $U_{\alpha}\propto\mathcal{V}(2k_{F})$, $V_{\alpha}\propto\mathcal{V}(4k_{F})$ and they depend linearly and quadratically respectively, on the band cutoff $D$. In the following, we solve the interaction Hamiltonian (2) exactly, and treat the nonlinearities (4) perturbatively in small backscattering amplitudes [8].
@@ III Exact account of the interactions
In the absence of backscattering events, the continuous flow of chiral fermions along the QH edges, described by the Hamiltonian $H_{\rm p}=H_{0}+H_{C}$, represents the plasmonic excitations propagating along the chiral edges. These excitations can be accounted for using the standard equation-of-motion approach [36, 45, 46]. This allows one to express all bosonic fields $\Phi_{\alpha,\rm R/L}(x,\omega)$ in terms of the corresponding free fields $\Phi^{0}_{\alpha,\rm R/L}(x,\omega)$, which satisfy the well-known commutation relations and possess standard correlators
@@
To study transport properties, it is more convenient to introduce the compact bosonic fields $\Phi_{\alpha}\equiv\Phi_{\alpha,\rm R}-\Phi_{\alpha,\rm L}$, along with the corresponding free fields $\Phi^{0}_{\alpha}$. It is straightforward to show, by solving the equation of motion for the Hamiltonian $H^{\prime}\equiv H_{0}+H_{C}$, that the matrix $\Phi=\left(\Phi_{1},\Phi_{2},\Phi_{3},\Phi_{4},\Phi_{5}\right)^{\top}$ satisfies
@@
for some matrix $\mathbb{A}$ that depends on the interaction strength and number of edge channels. The matrix $\mathbb{N}$ accounting for the gate voltages takes the form
@@
We note that in Eq. (26), and in the following discussion, we restrict ourselves to the regime $T\ll E_{c}$, where the physics is dominated by quantum fluctuations. In typical experiments, such as those in Refs. [10, 14], the base temperature is $T=(10-20)\,\rm{mK}$, and the charging energy is $E_{c}=(25-30)\,\mu\rm{eV}\sim(290-348)\,\rm{mK}$, which is well within the regime of our interest. The opposite regime, $T\gg E_{c}$, can be investigated in the spirit of Ref. [44]. From Eqs. (25) and (26), it follows that out of five chiral bosonic modes, only one mode, $\Phi_{a}$, remains gapless, while the remaining four acquire gap due to the large charging energy. For the low-energy description, the gapped modes can be integrated out, leaving the dynamics governed solely by the free mode $\Phi_{a}$. Equation (26) further implies that the scaling dimensions of the boundary perturbations appearing in the first and second terms of the equation (4) are $1/5$ and $4/5$ respectively, and therefore relevant perturbations in RG sense.